OpenStudy (anonymous):

simplify 2xv^7 . 6w^-2 . 5x^8 w^3 v^-9

OpenStudy (anonymous):

mathway.com

OpenStudy (radar):

First multiply the coefficients 2, 6, 5, This gives you 60. next you multiply the variables using the rule that multiplication is accomplished by adding the exponents.\[x \times x ^{8}= x ^{9} \]Now do the next variable\[v ^{7}\times v ^{-9}= v ^{-2}\] contoinueing with the next variable\[w ^{-2}\times w ^{3}=w\]Now put it all together.\[60x ^{9}wv ^{-2}\]

OpenStudy (anonymous):

wrong answer

OpenStudy (anonymous):

@radar